Find the fundamental price in one year of P1. The annual dividend of Safari Incorporated in the coming year is expected to be D1 = $2.25 and management expects to maintain a dividend growth rate of g = 3% indefinitely. Investors require a return of R i = 10% on this type of stock. What is this stock's expected price in one year or P1?

29.89

32.65

$31.75

$33.11

$29.20
Answer #1

Answer: $33.11

Explanation:

Stock Expected price in one year P1

= D2/(k-g)

= (2.25*103%)/(10%-3%) = $33.11